Accelerator Update for January 23, 2009 - January 26, 2009 |
Friday Jan 23 The day shift began with Operations monitoring the Tevatron (TeV) store 6748 (55.79E30), the Recycler stash (352.1E10), and the Antiproton Source stack (26.9E10), with the Linac and Booster operating normally, with Main Injector (MI) suffering from problems with three kickers (I:KPS4A & B and I:KPS5S), a sextupole power supply (SEXHPS), and with clock (TCLK) repeater problems, with the MiniBooNE experiment taking beam, but with NuMI only taking beam intermittently due to MI problems and their far detector being calibrated, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. Controls experts reported at 8:07 AM that they were working on the TCLK repeaters and ramp cards at MI-30 and 40. At 8:07 AM, a Booster RF station (BRF8) tripped off for the first of two times in 19 minutes. It reset. The MI ramp was reestablished at 8:42 AM. Operations established beam to NuMI at 10:38 AM. Pbar resumed stacking at 11:23 AM. Operations terminated store 6748 at 12:35 PM. RF experts began cleaning the TeV RF LCW strainers at 12:47 PM. At 1:17 PM, SEXHPS tripped off. Experts investigated and had the power supply back up by 2:35 PM. Operations put the TeV into shot setup at 2:41 PM. Operations established store 6749 at 4:31 PM with an initial luminosity of 290.5E30. At 4:44 PM, Operations turned off MI to reset some control cards (453 cards) in various CAMAC crates. Pbar resumed stacking and NuMI resumed taking beam at 5:01 PM. Operations used protons to check out the Recycler at 5:12 PM. Everything looked good. At 9:04 PM, MIRF8 tripped off. Operators investigated and found that an internal fan had failed. They bypassed the station and reconfigured MIRF6 to work as a slipstacking station to take its place. Operations reported at midnight that the Recycler had stashed antiprotons seven times during the evening shift. Saturday Jan 24 The midnight shift began Operations monitoring store 6749 (110.32E30) and the stash (217.0E10), with Pbar stacking (33.7E10), and with NuMI and MiniBooNE taking beam. Operations terminated store 6749 at 7:23 AM. Operations put the TeV into shot setup at 7:35 AM. Operations reported at 8:04 AM that the Recycler had stashed nine times during the midnight shift. Operations established store 6750 at 9:46 AM with an initial luminosity of 286.93E30. At 10:05 AM, CDF reported having problems with their plug high voltage. At 10:27 AM, Operations reported that the Tevatron Electron Lens (TEL2) was not coming up to its proper current. Operations contacted an expert who had it working normally by 11 AM. At 1:11 PM, a Linac Klystron RF station (KRF5) tripped off. It reset. At 6:20 PM, Operations reported having a problem with stashing antiprotons into the Recycler. One transfer was only 60% efficient and the next never made it into MI due to a kicker (I:KPS62A) trip. The kicker reset. At 7:55 PM, Operations reported that the MI-20 lower power supply bypass interlock loop had tripped off twice. Operations contacted experts. Operations put Pbar into standby mode at 7:57 PM. Operations reported at 8 PM that the Recycler had stashed 11 times during the last 12 hours. At 10:51 PM, all the high-energy Linac Klystron RF stations tripped off on flow alarms. Operators investigated and discovered that water skid #3 had suffered a pressure drop. It reset. At 11:42 PM, Operations reported that column #1 at CUB suffered a pressure drop. At 11:44 PM, Operations reported that the MI LCW system tripped off. This caused the Recycler correction elements to trip off: the stash was lost. At 11:51 PM, operators discovered a MI LCW leak at MI-20. One of the LCW pumps blew a gasket. Operations contacted the Duty Mechanic and called in water experts.
Sunday Jan 25 The midnight shift began with Operations monitoring store 6750, the stash, and the Pbar stack, with MI off due to an LCW problem, and with NuMI and MiniBooNE taking beam. At 12:01 AM, the Linac Klystron water skid #3 tripped off again. It reset. At 12:08 AM, the Booster East Anode Power Supply tripped off and would not reset. Operations conferred with an expert and brought it back online. At 12:10 AM, Operations reported that the Pelletron was tripping off and then would not reset. At 3:16 AM, the Duty Mechanic reported that when they started the MI-20 LCW pump it failed. The Duty Electrician investigated. Operations established beam to MiniBooNE at 3:22 AM. Experts reported that the LCW pumps at MI-20 were fixed. Operations resumed sending beam to Pbar and NuMI at 5:37 AM. At 7:10 AM, KRF3 tripped off for the first of two times in about 30 minutes. It reset. Operations reported at 8:01 AM that the Recycler had stashed five times during the last 12 hours. Operations terminated store 6750 at 12:31 PM. Operations put the TeV into shot setup at 12:50 PM. At 2:05 PM, BRF3 tripped off for the first of seven times in 14 minutes. It reset. At 2:33 PM, the Booster West Anode Power Supply and BRF13 tripped off. Neither of them would reset. Operations called in an expert. Operations established store 6751 at 2:48 PM with an initial luminosity of 159.06E30. An RF expert reported at 4:17 PM that he found and repaired a failed fan in BRF13. Operations resumed sending beam to NuMI and Pbar. Operations reported at 7:59 PM that the Recycler had stashed nine times during the last 12 hours. Monday Jan 26 The midnight shift began with Operations monitoring store 6751 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. At 4:08 AM, the 750KeV area buncher tripped off. It reset. Operations terminated store 6751 at 7:24 AM. Operations put the TeV into shot setup at 7:39 AM. Operations reported at 8:03 AM that the Recycler had stashed 14 times during the last 12 hours. The Plans for Monday The plans for today are to allow D0 a two-hour access to fix a power supply and then put in a new store later in the morning. |
|
More Information Comments and Suggestions |
| last modified 01/26/2009 email Fermilab |
| Security, Privacy, Legal | |